OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form.

OpenStudy (curry):

-x + y = 7

OpenStudy (anonymous):

Use this equation: \[y - y_{1} = m(x- x_{1})\] with y1 = 5, x1 = -1, and m = 2

OpenStudy (anonymous):

y - 5 = 2(x -(-1)) y - 5 = 2(x + 1) y - 5 = 2x + 2 y = 2x + 7